OpenStudy (anonymous):

How do I solve for t in the equation \[\frac{t(t^2-2)}{t^2-1}=0\] I know the answers are +-square root of 2 and 0 but that's after I punch it into my calculator. Could you show me the steps if I want to evaluate this by hand? Thanks!

OpenStudy (anonymous):

Since the equation is equal to zero, only the numerator is important.

OpenStudy (anonymous):

Why only the numerator is important?

OpenStudy (anonymous):

If t=0, the entire equation will equal 0, likewise, if t^2-2=0 the entire equation will equal 0.

jimthompson5910 (jim_thompson5910):

If A/B = 0, then multiply both sides by B to get A = B*0 ---> A = 0

jimthompson5910 (jim_thompson5910):

that's why only the numerator matters

jimthompson5910 (jim_thompson5910):

besides, you can't divide by zero

OpenStudy (anonymous):

don't I have to check for \[t^2-1 \neq 0\]?

jimthompson5910 (jim_thompson5910):

yes it's best to find out which numbers make the denominator zero first

OpenStudy (anonymous):

That would give t=1 or t=-1. Then the solution to t would t=+-sqrt(2), t=0 and t not equals +-1?

jimthompson5910 (jim_thompson5910):

With the solution, you only need to worry about what t equals (instead of what it can't equal)

OpenStudy (anonymous):

Hey jim, sorry to butt in on this conversation, but do you know how to create a fraction in the equation editor? It is very difficult to give coherent answers without fractions.

jimthompson5910 (jim_thompson5910):

type \frac{x}{y} to type out \[\Large \frac{x}{y}\]

jimthompson5910 (jim_thompson5910):

for some reason, this is missing in the equation editor...

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Suppose another function call it E(t) holds only if the above equation of \[\frac{t(t^2-2)}{t^2-1} \neq 0\], Then, for E(t) to hold, t must be \[t \neq \pm \sqrt{2}\] \[t \neq 0\] But then what if the function was given -1 or 1 into it? The whole thing would not work since it will be division by zero, wouldn't it?

OpenStudy (anonymous):

As in, since the condition that I have got from the equation, was just \[t \neq \pm \sqrt{2} \; \; \; and \; \; \; t \neq 0\] Then wouldn't there be a possibility/danger that someone throws in a value of t which is -1 or 1?

jimthompson5910 (jim_thompson5910):

yes that possibility exists, but you just ignore those potential solutions if it arises

OpenStudy (anonymous):

Correct, the function would not work if \[t=\pm1\], one way to avoid this would be to define the domain, so t=\[(-\infty,-1),(-1,1),(1,\infty)\]

OpenStudy (anonymous):

Thanks everyone! :)

OpenStudy (anonymous):

You're welcome.